Two identical containers joined by a small pipe initially contain the same gas at pressure Po and absolute temperature To. One container is now contained at the same temperature while the other is heated to 2To. Find the common pressure of the gases.

Grade:12

2 Answers

ankitesh gupta
11 years ago
ANSWER IS 2PO/3
IN THIS QUESTION YOU HAVE TO APPLY CONSERVATION OF MOLES
FIRST CONSIDER TWO BLOCKS OF VOLUME ''V'' PRESSURE ''P0'' AND TEMPRATURE '' TO'' THEREFORE TOTAL MOLES IN THE TWO BLOCKS IS
PO (2V) = nRTO     ....(1)
NOW LET THE COMMON PRESSURE LATER BE ''P'' AND CONSIDER ONE BLOCK TO HAVE TEMPRATURE  TO  AND NO. OF MOLES TO BE n1 AND THE OTHER BLOCK WILL HAVE TEMPRATURE 2TO AND NO. OF MOLES TO BE n2 THEN  
FOR 1ST BLOCK..... PV = n1RTO         .....(2)
FOR 2ND BLOCK .....PV = n2R(2TO)      .....(3)
NOW SINCE TOTAL MOLES ARE CONSERVED
n1+n2 = n                             .....(4)
FROM EQUATION (1) , (2) ,(3), (4)
WE GET P=2PO/3 ANSWER
PLZ APPROVE THE ANSWER IF YOU ARE SATISFIED BY CLICKING THE YES BUTTON
San jeep kumar
6 years ago
Answer is 4p○/3 not 2p○/3 IN THE ABOVE ANSWER EVERYTHING IS CORRECT JUST CALCULATION MISTAKE after n1+n2=n we get 2p○= p+p/2 so which result in 4p○ /3